# Central and Central-Parietal EEG Signatures of Parkinson’s Disease

### v1

@@ 4.1.2 Central and Parietal Involvement: Neuroanatomical Significance
The superior performance of central electrodes (76% accuracy with 20-second windows) directly corresponds to PD’s impact on motor circuitry. Similarly, the high classification accuracy in bilateral parietal regions (73%) aligns with Borghammer’s findings of reduced parietal blood flow in PD patients [19]. These parietal abnormalities may contribute to deficits in spatial awareness and sensorimotor integration evident even in early disease stages [20].
@@ 4.1.3 Frequency-Specific Alterations
The exceptionally high gamma band (40-62.4 Hz) accuracy in central-parietal electrodes (74%) represents our most striking finding. This suggests PD significantly disrupts high-frequency synchronization in regions critical for sensorimotor integration. While gamma oscillations typically facilitate feature binding and attention [22], their specific alteration in PD has been underexplored. These midline regions integrate diverse sensory inputs for coordinated movement, with gamma oscillations potentially binding this information across cortical areas. The observed gamma alterations may contribute to the movement initiation and coordination difficulties characteristic of PD.
Our finding of high gamma band discriminative power aligns with the emerging understanding of gamma oscillations in PD. Two distinct types of gamma activity in PD were identified: a ”finely-tuned gamma” (70-80 Hz) observed only in the ON medication state and typically associated with dyskinesias, and a broader ”movement-related” gamma that increases during voluntary movement [23]. The gamma band (40-62.4 Hz) alterations we observed in the central-parietal and right parieto-occipital regions might represent a resting-state manifestation of abnormal gamma processing that becomes more pronounced during movement. Importantly, our findings extend previous knowledge by demonstrating that these gamma alterations can be detected in scalp EEG recordings, not just in invasive local field potential (LFP) recordings from deep brain structures as most previous research has shown.
The bilateral centro-parietal theta band (4.0-7.79 Hz) alterations (67% accuracy) may partially reflect tremor-related activity that remains after artifact cleaning, as parkinsonian tremor typically falls within this frequency range. Additionally, these alterations likely reflect disruptions in cognitive processes including working memory and attention—functions frequently compromised in PD [6]. Prior research has documented increased theta activity in PD [24], possibly reflecting compensatory mechanisms or pathological synchronization.
The widespread alpha band (7.8-15.59 Hz) alterations across multiple regions align with extensive literature on alpha rhythm disruption in PD [25, 26]. These changes potentially indicate dysfunction in thalamocortical circuits that normally generate and maintain alpha rhythms [27].
@@ 4.2 Neurophysiological Mechanisms
Recent comprehensive reviews of brain oscillations in PD conceptualize the disease as an ”oscillopathy” [23], with distinct oscillatory patterns associated with different clinical states and symptoms. Within this framework, our findings of altered gamma and theta oscillations in specific cortical regions contribute to this broader understanding of PD’s oscillatory signatures.
Several complementary mechanisms likely underlie our observed EEG alterations:
@@ 4.2.1 Dopaminergic Pathways and Network-Wide Effects
PD’s characteristic dopaminergic depletion disrupts the balance between facilitatory and inhibitory basal ganglia pathways, ultimately altering thalamocortical circuits [1]. This disruption manifests in our results as altered oscillatory patterns, particularly in frontal/fronto-central regions most directly connected to these pathways. The relationship between dopamine levels and EEG features is supported by George et al.’s demonstration that dopaminergic medication decreases cortical beta coherence in PD patients [15].
These primary pathological changes interact with compensatory mechanisms, creating what Voytek and Knight termed ”network dysrhythmia” [16]—abnormal oscillatory patterns across interconnected neural networks. This framework helps explain why different brain regions showed distinct frequency preferences in our analysis. The central-parietal regions’ dual sensitivity in both theta and gamma bands, contrasting with other regions’ different frequency preferences, suggests PD affects various functional networks through distinct mechanisms, potentially involving multiple neurotransmitter systems.
@@ 4.2.2 Sensorimotor Integration and Network Connectivity
The exceptional gamma band discriminative power in central-parietal regions likely represents disrupted sensorimotor integration processes. These midline areas integrate multiple sensory inputs crucial for coordinated movement. Recent evidence indicates that pathological $\alpha$-synuclein aggregation affects not only basal ganglia but multiple brain regions through various mechanisms [28], potentially explaining these cortical oscillatory alterations.
The alpha-band sensitivity in longitudinal connections further supports a network-level disruption model. Alpha oscillations typically coordinate activity across large-scale brain networks, and their disruption may reflect impaired communication between frontal control regions and posterior sensory areas—potentially contributing to the sensorimotor integration deficits characteristic of PD.
@@ 4.2.3 Cross-frequency Coupling and Network Dynamics
While our study focused on power changes across different frequency bands, recent work has highlighted the importance of cross-frequency coupling in PD pathophysiology. In particular, as mentioned earlier Foffani and Alegre describe significant coupling between beta oscillations and high-frequency oscillations in the basal ganglia of PD patients, with the strength of this coupling correlating with symptom severity [23]. Future EEG studies should investigate whether such cross-frequency coupling can be detected at the scalp level, potentially providing even more sensitive biomarkers than power changes in individual frequency bands. Our finding of concurrent alterations in multiple frequency bands (gamma, theta, and alpha) across different but overlapping cortical regions suggests possible disruptions in the normal coupling relationships between these oscillations, which may reflect fundamental network-level dysfunction in PD.
@@ 4.3 Contributions and Implications
Our results identify specific biomarker candidates, particularly central-parietal gamma activity and central region full-spectrum activity, which achieved the highest classification accuracies. These spatial-spectral patterns could potentially be developed into targeted biomarkers for PD detection.
@@
The aforementioned findings is especially important in the context of beta oscillations and their coupling with high-frequency oscillations, as they are increasingly being used as control signals for adaptive deep brain stimulation [23]. While these approaches currently rely on invasive recordings, our demonstration that discriminative oscillatory patterns can be detected non-invasively through EEG suggests potential pathways toward less invasive neuromodulation approaches. The spatial specificity of our findings—particularly the central-parietal gamma and right parieto-occipital gamma patterns—might inform the development of targeted non-invasive stimulation protocols using techniques such as transcranial alternating current stimulation at these specific frequencies and locations.
@@ 4.4.1 Segment-Level Classification and Variability
Despite this window-level variability, the consistency of spatial-spectral patterns across different window lengths supports our findings’ robustness. The superior performance of longer windows (20 seconds) compared to shorter ones (5 or 10 seconds) suggests PD-related EEG alterations manifest as subtle but persistent oscillatory pattern changes rather than transient events.
@@ 4.4.2 Classification Accuracy in Context
5 Limitations and Future Directions
@@ 5.0.3 Future Research Directions
The spatial-spectral EEG patterns identified in this study—particularly the central-parietal gamma activity and centro-parietal theta alterations—suggest several targeted avenues for future investigation:
Most critically, our findings of 74% classification accuracy in the gamma band for right parieto-occipital electrodes (P4, O4, PO4) warrant validation in prodromal PD populations. Specifically, longitudinal monitoring of individuals with idiopathic REM sleep behavior disorder (iRBD), who convert to clinical PD at rates of 40-75% within 10 years, could determine whether these gamma-band abnormalities appear before motor symptoms emerge.
The pronounced right-hemisphere involvement identified across several frequency bands raises important questions about lateralization of early pathophysiological changes in PD. Future research should examine whether this asymmetry correlates with side of initial motor symptom presentation, which would link our EEG findings directly to the known asymmetric degeneration pattern in early PD. The high discriminative power in gamma bands (40-62.4 Hz) specifically calls for experimental designs incorporating tasks known to modulate gamma activity, such as visual processing paradigms, to determine whether task-evoked gamma responses might provide even stronger classification accuracy than resting-state measures.
@@ 6 Conclusion
This study mapped PD-related EEG alterations across multiple cortical regions and frequency bands, revealing specific spatial-spectral signatures with diagnostic potential. Most notably, we identified a pronounced central-parietal gamma band (40-62.4 Hz) alteration that achieved 74% classification accuracy, alongside robust discriminative power (76%) in central electrodes using full-spectrum analysis. Unexpectedly, right hemisphere involvement was prominent, particularly in parieto-occipital regions, challenging the traditional focus on left-dominant manifestations in PD literature.
The remarkable performance of gamma oscillations in sensorimotor integration regions directly links our findings to emerging theories of network dysrhythmia in PD. The 20-second window’s superior performance over shorter segments suggests these oscillatory abnormalities represent stable pathophysiological features rather than transient phenomena. The observed theta band sensitivity in bilateral centro-parietal regions and widespread alpha alterations further substantiate a complex, region-specific pattern of neural synchronization disruption that extends beyond the classical cortico-basal ganglia motor circuits.
Future longitudinal studies examining these markers in prodromal populations will be essential to support our results and employ them for early detection and intervention in Parkinson’s disease.

### v2

@@ 4.1.2 Central and Parietal Involvement: Neuroanatomical Significance
The superior performance of central electrodes (76% accuracy with 20-second windows) directly corresponds to PD’s impact on motor circuitry. Similarly, the high classification accuracy in right parietal regions (74% with 10-second windows) aligns with Borghammer’s findings of altered parietal function in PD patients [21]. These parietal abnormalities may contribute to deficits in spatial awareness and sensorimotor integration evident even in early disease stages [22].
The asymmetric involvement, with right parietal regions showing stronger discriminative power than their left counterparts at shorter window lengths, merits particular attention. This right-hemisphere emphasis challenges the traditional view that PD-related cortical changes mirror the often asymmetric motor symptoms that frequently begin on the left side. Several functional neuroimaging studies have similarly reported right-hemisphere vulnerability in PD, particularly for non-motor functions [14]. This lateralization may reflect the right hemisphere’s specialized role in spatial attention and sensory integration, functions known to be affected in PD.
@@ 4.1.3 Frequency-Specific Alterations
The bilateral centro-parietal theta band (4.0-7.79 Hz) alterations (67% accuracy) may reflect changes in cognitive processes including working memory and attention—functions frequently compromised in PD [6]. Prior research has documented increased theta activity in PD [24], possibly reflecting compensatory mechanisms or pathological synchronization in cortical-subcortical circuits. The specificity of theta band sensitivity to bilateral centro-parietal regions is particularly noteworthy, as these regions are known to be involved in sensorimotor integration functions.
The widespread alpha band (7.8-15.59 Hz) alterations across multiple regions align with extensive literature on alpha rhythm disruption in PD [25, 26]. Several regions achieved approximately 65% accuracy in the alpha band, including central, right parietal, and longitudinal connections. This broad spatial distribution of alpha alterations suggests a more global impact on thalamocortical circuits that normally generate and maintain alpha rhythms [14]. The ubiquity of alpha band alterations across functionally distinct regions may reflect the widespread network effects of PD pathophysiology.
The left centro-parietal beta1 band (15.6-31.19 Hz) performance (65% $\pm$ 31%) is particularly interesting given the known importance of beta oscillations in PD pathophysiology. Beta oscillations typically reflect motor preparation and inhibition processes, and their disruption may contribute to the movement initiation difficulties characteristic of PD [20]. The localization of beta1 alterations to left centro-parietal regions may reflect hemisphere-specific changes in motor control networks.
While individual frequency bands showed region-specific sensitivity, the highest overall classification accuracies were achieved with full-spectrum analysis (76% for central electrodes, 74% for right parietal), suggesting that the complex patterns of oscillatory changes across multiple frequency bands collectively provide the most discriminative information. This multi-spectral signature likely better captures the complex dysrhythmia that characterizes PD than any single frequency band alone.
@@ 4.2 Neurophysiological Mechanisms
Recent comprehensive reviews of brain oscillations in PD conceptualize the disease as an ”oscillopathy” [20], with distinct oscillatory patterns associated with different clinical states and symptoms. Within this framework, our findings of altered theta, alpha, and beta oscillations in specific cortical regions contribute to this broader understanding of PD’s oscillatory signatures.
4.2.1 Network Dysrhythmia Framework
The ”Network Dysrhythmia Framework” proposed by Voytek and Knight [17] provides a lens for interpreting our findings, conceptualizing neurological disorders as disruptions in rhythmic activity patterns across interconnected neural networks. Our results align with this framework in several important ways.
Our observation of frequency-region specificity—such as theta dominance in centro-parietal regions, widespread alpha alterations, and focused beta1 changes in left centro-parietal areas—suggests PD creates distinct dysrhythmic patterns in different functional networks rather than causing global oscillatory changes. These region-specific alterations likely reflect differential vulnerability to PD pathology or engagement in compensatory processes among neural circuits.
The disruptions we identified operate across multiple spatial scales, ranging from localized electrode triplet effects to broader hemisphere-level asymmetries. This multi-level impact on neural synchronization affects both local processing and long-range communication, with the right-hemisphere emphasis underscoring PD’s spatially heterogeneous effects on brain networks.
Additionally, the persistence of these discriminative oscillatory features across different window lengths (5, 10, and 20 seconds) indicates stable alterations in brain dynamics rather than transient phenomena, consistent with PD’s chronic, progressive impact on neural circuits.
This network perspective moves beyond viewing PD solely as a dopaminergic deficit in the basal ganglia and instead recognizes it as a systems-level disorder affecting synchronization patterns across distributed cortical networks. Such a framework better explains both motor and non-motor symptoms and the regional specificity of oscillatory alterations we observed.
@@ 4.2.2 Dopaminergic Pathways and Network-Wide Effects
PD’s characteristic dopaminergic depletion disrupts the balance between facilitatory and inhibitory basal ganglia pathways, ultimately altering thalamocortical circuits [1]. This disruption manifests in our results as altered oscillatory patterns, particularly in central regions most directly connected to these pathways. The relationship between dopamine levels and EEG features is supported by George et al.’s demonstration that dopaminergic medication decreases cortical beta coherence in PD patients [16].
These primary pathological changes interact with compensatory mechanisms, creating abnormal oscillatory patterns across interconnected neural networks. This framework helps explain why different brain regions showed distinct frequency preferences in our analysis. The central-parietal regions’ sensitivity in the theta band, contrasting with other regions’ different frequency preferences, suggests PD affects various functional networks through distinct mechanisms, potentially involving multiple neurotransmitter systems.
@@ 4.2.3 Sensorimotor Integration and Network Connectivity
The strong discriminative power in both central and parietal regions likely represents disrupted sensorimotor integration processes. These areas integrate multiple sensory inputs crucial for coordinated movement. Recent evidence indicates that pathological $\alpha$-synuclein aggregation affects not only basal ganglia but multiple brain regions through various mechanisms [27], potentially explaining these cortical oscillatory alterations.
The alpha-band sensitivity in longitudinal connections and parieto-occipital regions further supports a network-level disruption model. Alpha oscillations typically coordinate activity across large-scale brain networks, and their disruption may reflect impaired communication between frontal control regions and posterior sensory areas—potentially contributing to the sensorimotor integration deficits characteristic of PD.
@@ 4.2.4 Cross-frequency Coupling and Network Dynamics
While our study focused on power changes across different frequency bands, recent work has highlighted the importance of cross-frequency coupling in PD pathophysiology. In particular, Foffani and Alegre describe significant coupling between beta oscillations and high-frequency oscillations in the basal ganglia of PD patients, with the strength of this coupling correlating with symptom severity [20].
Our finding of concurrent alterations in multiple frequency bands (theta, alpha, and beta1) across different but overlapping cortical regions suggests possible disruptions in the normal coupling relationships between these oscillations, which may reflect fundamental network-level dysfunction in PD. Future EEG studies should investigate whether such cross-frequency coupling can be detected at the scalp level, potentially providing even more sensitive biomarkers than power changes in individual frequency bands.
@@ 4.3 Contributions and Implications
Our results identify specific biomarker candidates, particularly central region full-spectrum activity (76% accuracy) and right parietal region activity (74% accuracy), which achieved the highest classification accuracies. The strong theta band performance in bilateral centro-parietal regions and widespread alpha band sensitivity also represent potential targeted biomarkers for PD detection.
By systematically analyzing different electrode combinations across the scalp, we identified the involvement of parietal regions, particularly right parietal areas, which have received less attention than frontal and central regions in previous PD EEG research. This highlights the importance of comprehensive spatial sampling rather than focusing solely on regions with known motor involvement.
Our analysis across different window lengths (5, 10, and 20 seconds) provides insights into the temporal characteristics of discriminative EEG features in PD. The minimal differences in overall mean accuracy across window lengths (58.4% to 60.2%) suggest that PD-related EEG alterations are detectable across various time scales, though specific regions like central motor areas may benefit from longer sampling periods. This finding has practical implications for clinical applications, suggesting that even relatively brief EEG recordings may contain diagnostically useful information.
@@
These findings are especially important in the context of oscillatory activity being increasingly used as control signals for adaptive deep brain stimulation [20]. While these approaches currently rely on invasive recordings, our demonstration that discriminative oscillatory patterns can be detected non-invasively through EEG suggests potential pathways toward less invasive neuromodulation approaches. The spatial specificity of our findings—particularly the central and right parietal patterns—might inform the development of targeted non-invasive stimulation protocols using techniques such as transcranial alternating current stimulation at specific frequencies and locations.
@@ 4.4.1 Segment-Level Classification and Variability
The minimal differences in mean classification accuracy across window lengths (58.4% to 60.2%) suggest that PD-related EEG alterations are detectable at various temporal scales, though region-specific effects of window length were observed, particularly for central electrodes which showed substantial improvement with longer windows (68% to 76% from 10 to 20 seconds).
@@ 4.4.2 Classification Accuracy in Context
4.5 Limitations and Future Directions
@@ 4.5.3 Future Research Directions
The spatial-spectral EEG patterns identified in this study suggest several targeted avenues for future investigation:
Most critically, our findings of high classification accuracy in the central and right parietal regions warrant validation in prodromal PD populations. Specifically, longitudinal monitoring of individuals with idiopathic REM sleep behavior disorder (iRBD), who convert to clinical PD at rates of 40-75% within 10 years, could determine whether these oscillatory abnormalities appear before motor symptoms emerge.
The pronounced right-hemisphere involvement identified across several frequency bands raises important questions about lateralization of early pathophysiological changes in PD. Future research should examine whether this asymmetry correlates with side of initial motor symptom presentation, which would link our EEG findings directly to the known asymmetric degeneration pattern in early PD.
The strong discriminative power in theta bands specifically calls for experimental designs incorporating tasks known to modulate theta activity, such as working memory paradigms, to determine whether task-evoked theta responses might provide even stronger classification accuracy than resting-state measures. Similarly, the widespread alpha band alterations warrant investigation using paradigms that typically engage alpha oscillations, such as attentional tasks and visual processing.
Future studies should also investigate cross-frequency coupling patterns, particularly between theta-alpha and alpha-beta frequencies, which may provide more sensitive markers of PD-related network dysfunction than power in individual frequency bands. Recent advances in connectivity analysis methods could further illuminate how PD disrupts communication between brain regions at different temporal scales.
@@ 5 Conclusion
The remarkable performance of central and parietal regions directly links our findings to theories of network dysrhythmia in PD. The minimal differences in mean accuracy across window lengths suggests these oscillatory abnormalities are detectable at various temporal scales, indicating PD-related EEG features have temporal consistency. Individual electrode triplets like C3, CZ, C4 did show more substantial improvements with longer windows (68% to 76% from 10 to 20 seconds).
The observed theta band sensitivity in bilateral centro-parietal regions, widespread alpha alterations, and spatially-specific gamma patterns substantiate a complex, region-specific pattern of neural synchronization disruption that extends beyond the classical cortico-basal ganglia motor circuits.
Our methodology, which carefully removed tremor-related artifacts, ensured that classification was based on genuine neurophysiological alterations rather than mechanical tremor detection. This approach, though yielding more modest accuracy figures than some previous reports, provides greater confidence that the identified biomarkers reflect intrinsic brain oscillatory changes associated with PD pathophysiology.
Future longitudinal studies examining these markers in prodromal populations will be essential to determine their utility for early detection and intervention in Parkinson’s disease.